# Nearsightedness in Materials with Indirect Band Gap

### v1

@@ Appendix B Computational Details in 1D Chain
Here we will provide the computational details of 1D chain tests. Applying Bloch transform for the Hamiltonian in (4.1) yields that
${\mathcal{H}}_{\xi}=\left(\begin{array}[]{cc}c_{1}+2f_{1}(1)\cos(\xi)&2f_{3}(1)\cos(\xi)\\ 2f_{3}(1)\cos(\xi)&c_{2}+2f_{2}(1)\cos(\xi)\end{array}\right),\qquad\forall\xi\in\Gamma^{*}=[-\pi,\pi].$
For the eigenvalue problem ${\mathcal{H}}_{\xi}u_{n}(\xi)=\varepsilon_{n}(\xi)u_{n}(\xi)$, we solve it by considering the equation
$\det\big{|}\varepsilon(\xi)-{\mathcal{H}}_{\xi}\big{|}=0,$
we obtain the two energy bands $\{\varepsilon_{n}(\xi)\}_{n=1,2}$,
$\begin{split}\varepsilon_{1}(\xi)&=\frac{a(\xi)+b(\xi)}{2}-\sqrt{\frac{(a(\xi)-b(\xi))^{2}}{4}+4\left(f_{3}(1)\cos(\xi)\right)^{2}}\\ \varepsilon_{2}(\xi)&=\frac{a(\xi)+b(\xi)}{2}+\sqrt{\frac{(a(\xi)-b(\xi))^{2}}{4}+4\left(f_{3}(1)\cos(\xi)\right)^{2}}\end{split}$
where we denote $a(\xi):=c_{1}+2f_{1}(1)\cos(\xi)$ and $b(\xi):=c_{2}+2f_{2}(1)\cos(\xi)$. Then $\mathsf{gap}_{+}$ can be given immediately by
$\begin{split}\mathsf{gap}_{+}&=\min_{\xi\in\Gamma^{*}}\big{(}\varepsilon_{2}(\xi)-\varepsilon_{1}(\xi)\big{)}=\min_{\xi\in\Gamma^{*}}2\sqrt{\frac{(a(\xi)-b(\xi))^{2}}{4}+4\left(f_{3}(1)\cos(\xi)\right)^{2}}\\ &=\frac{2f_{3}(1)\left|c_{1}-c_{2}\right|}{\sqrt{\big{(}f_{1}(1)-f_{2}(1)\big{)}^{2}+\big{(}2f_{3}(1)\big{)}^{2}}}.\end{split}$ (B.1)
By adjusting $f_{1}(1),f_{2}(1),f_{3}(1),c_{1},c_{2}$, we can get systems with fixed $\mathsf{gap}_{+}$ but different $\mathsf{gap}_{-}$.
Remark B.1. There exits other possible expression for $\mathsf{gap}_{+}$:
$\mathsf{gap}_{+}=\big{[}c_{1}-c_{2}\pm 2(f_{1}(1)-f_{2}(1))\big{]}^{2}+16f_{3}(1)^{2},$
as the minima of $\big{(}\varepsilon_{2}(\xi)-\varepsilon_{1}(\xi)\big{)}$ occur at 00 or ±$\pm\pi$. However, in our tests, the chosen parameters yield $\mathsf{gap}_{+}$ as defined by (B.1). ∎
@@ Appendix C Exponential Sums
$\sum_{m\in A}e^{-\gamma r_{\ell m}}e^{-\eta r_{mk}}$ $\leq\left[1+\frac{C_{d}}{(\mathfrak{m}\gamma)^{d}}\right]e^{-\frac{1}{2}\min\limits_{m\in A}[\gamma r_{\ell m}+\eta r_{mk}]}$ (C.1)
$\leq\left[1+\frac{C_{d}}{(\mathfrak{m}\gamma)^{d}}\right]e^{\gamma\sup\limits_{a\in A}|a|}e^{-\frac{1}{2}(\gamma|\ell|+\eta|k|)}$ (C.2)
@@
$\sum_{m\in A}e^{-\gamma r_{\ell m}}e^{-\eta r_{mk}}$ $\leq\Big{(}\sum_{m\in A}e^{-\frac{1}{2}\gamma r_{\ell m}}\Big{)}e^{-\frac{1}{2}\min\limits_{m\in A}[\gamma r_{\ell m}+\eta r_{mk}]}.$ (C.3)
[…]
$\sum_{m\neq\ell}e^{-\frac{1}{2}\gamma r_{\ell m}}$ $\leq\sum_{m\neq\ell}\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.57222pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.70335pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.59167pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.10834pt}}\!\int_{B_{r_{\ell m}}(\bm{r}_{\ell})\cap B_{\mathfrak{m}/2}(\bm{r}_{m})}e^{-\frac{1}{2}\gamma|\bm{r}-\bm{r}_{\ell}|}{\,{\rm d}}\bm{r}\leq\frac{C_{d}}{\mathfrak{m}^{d}}\int_{\mathbb{R}^{d}}e^{-\frac{1}{2}\gamma|\bm{r}-\bm{r}_{\ell}|}{\,{\rm d}}\bm{r}$

### v2

@@ Appendix B Computational Details in 1D Chain
Here we will provide the computational details of 1D chain tests. The Bloch transform gives that
${\mathcal{H}}_{\xi}={\bm{c}}+2{\bm{f}}(1)\cos\xi,\quad\forall\xi\in\Gamma^{*}=[-\pi,\pi].$
We solve the eigenvalue problem of ${\mathcal{H}}_{\xi}$ by considering the equation $\det\big|\varepsilon(\xi)-{\mathcal{H}}_{\xi}\big|=0,$ which yields two energy bands denoted by $\varepsilon_{\pm}(\xi)$.
In particular, for the model of type (i) where $c_{12}=c_{21}=0$, we have
$\begin{split}\varepsilon_{\pm}(\xi)&=\frac{c_{11}+c_{22}}{2}+\big(f_{11}+f_{22}\big)\cos\xi\pm\sqrt{\frac{\big(c_{11}-c_{22}+2(f_{11}-f_{22}\cos\xi\big)^{2}}{4}+4\left(f_{12}\cos\xi\right)^{2}},\end{split}$
where we denote $f_{ij}\coloneqq f_{ij}(1)$ for $1\leq i,j\leq 2$. Then $\mathsf{gap}_{+}$ can be given immediately by
$\mathsf{gap}_{+}$ $=\min_{\xi\in\Gamma^{*}}\big(\varepsilon_{+}(\xi)-\varepsilon_{-}(\xi)\big)=\min_{\xi\in\Gamma^{*}}\sqrt{{\big(c_{11}-c_{22}+2(f_{11}-f_{22}\cos\xi)\big)^{2}}+16\left(f_{12}\cos\xi\right)^{2}}$
(B.1) $=\frac{2f_{12}\left|c_{11}-c_{22}\right|}{\sqrt{\big(f_{11}-f_{22}\big)^{2}+4f_{12}^{2}}}.$
Additionally, for the model of type (ii) where $f_{12}=f_{21}=0$, the associated energy bands are given by
$\begin{split}\varepsilon_{\pm}(\xi)=\frac{c_{11}+c_{22}}{2}+\big(f_{11}+f_{22}\big)\cos\xi\pm\sqrt{\frac{\big(c_{11}-c_{22}+2(f_{11}-f_{22}\cos\xi)\big)^{2}}{4}+4c_{12}^{2}},\end{split}$
where the $\mathsf{gap}_{+}$ is calculated by
$\mathsf{gap}_{+}=\min_{\xi\in\Gamma^{*}}\big(\varepsilon_{+}(\xi)-\varepsilon_{-}(\xi)\big)=\sqrt{\big(c_{11}-c_{22}\pm 2(f_{11}-f_{22})\big)^{2}+16c_{12}^{2}}.$
Next, we describe the parameters choices in our numerical experiments. For the type (i) model, we set $c_{11}=1+1/2\mathsf{gap}_{+},c_{22}=1-1/2\mathsf{gap}_{+}$ and
$f_{11}$ $=(0.5+0.5r)\eta(r),~f_{22}=(0.3+0.7r)\eta(r),$
$f_{13}$ $=(0.4+(p-0.4)r)\eta(r),$
where $p=\frac{1}{2}\sqrt{(\frac{1}{2}\mathsf{gap}_{-}+2)^{2}-\frac{1}{4}\mathsf{gap}_{+}}$ and $\eta(r)=e^{1-1/(1-|r-1|^{2})}$ as $r\in(1,2)$ which smoothly connects $\eta=1$ for $r\in[0,1]$ and $\eta=0$ for $r\in[2,\infty)$. Under this parametrization, the band functions take the form
$\varepsilon_{\pm}(\xi)=(1+2\cos\xi)\pm\frac{1}{2}\sqrt{\mathsf{gap}_{+}^{2}+[(\mathsf{gap}_{-}+4)^{2}-\mathsf{gap}_{+}^{2}]\cos^{2}\xi}.$
This allows us to design experiments with either fixed $\mathsf{gap}_{-}$ or fixed $\mathsf{gap}_{+}$. Moreover, as $\mathsf{gap}_{+}$ and $\mathsf{gap}_{-}$ tend to zero, we obtain the limiting band behaviour $\varepsilon_{\pm}(\xi)\approx 1+2(\cos\xi\pm|\cos\xi|)$ which corresponds to the case shown in Figure 4 .
For the second type, we set $c_{11}=-1,c_{22}=5,c_{12}=c_{21}=1/2\mathsf{gap}_{+},f_{11}(r)=(1+r)\eta(r)$ and $f_{22}(r)=(-2+r)\eta(r)$, which yields cosine-like bands
$\varepsilon_{\pm}(\xi)=(2+\cos\xi)\pm\frac{1}{2}\sqrt{\big(-6+6\cos\xi\big)^{2}+\mathsf{gap}_{+}^{2}}.$
As $\mathsf{gap}_{+}\to 0$, we observe that $\varepsilon_{+}(\xi)\approx-1+4\cos\xi$, and $\varepsilon_{-}(\xi)\approx 5-2\cos\xi$ which corresponds to the behavior illustrated in Figure 5.
Remark B.1. For the type (i), an alternative definition of $\mathsf{gap}_{+}$ could be used:
$\mathsf{gap}_{+}=\big[c_{11}-c_{22}\pm 2(f_{11}-f_{22})\big]^{2}+16f_{12}^{2},$
since the minima of $\varepsilon_{+}(\xi)-\varepsilon_{-}(\xi)$ occur at 0 or $\pm\pi$. However, in our tests, the chosen parameters yield $\mathsf{gap}_{+}$ as defined by (B). ∎
@@ Appendix C Exponential Sums
(C.1) $\sum_{m\in A}e^{-\gamma r_{\ell m}}e^{-\eta r_{mk}}$ $\leq\left[1+\frac{C_{d}}{(\mathfrak{m}\gamma)^{d}}\right]e^{-\frac{1}{2}\min\limits_{m\in A}[\gamma r_{\ell m}+\eta r_{mk}]}$
(C.2) $\leq\left[1+\frac{C_{d}}{(\mathfrak{m}\gamma)^{d}}\right]e^{\gamma\sup\limits_{a\in A}|a|}e^{-\frac{1}{2}(\gamma|\ell|+\eta|k|)}$
@@
(C.3) $\sum_{m\in A}e^{-\gamma r_{\ell m}}e^{-\eta r_{mk}}$ $\leq\Big(\sum_{m\in A}e^{-\frac{1}{2}\gamma r_{\ell m}}\Big)e^{-\frac{1}{2}\min\limits_{m\in A}[\gamma r_{\ell m}+\eta r_{mk}]}.$
Additionally, by noting that $e^{-\gamma r_{\ell m}}\leq e^{-\gamma|\bm{r}-\bm{r}_{\ell}|}$ for all $\bm{r}\in B_{r_{\ell m}}(\bm{r}_{\ell})$, and the non-interpenetration condition in (2.5), we can approximate eq. C.3 with the following integral
$\sum_{m\neq\ell}e^{-\frac{1}{2}\gamma r_{\ell m}}$ $\leq\sum_{m\neq\ell}\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.57222pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.91281pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.30664pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.13547pt}}\!\int_{B_{r_{\ell m}}(\bm{r}_{\ell})\cap B_{\mathfrak{m}/2}(\bm{r}_{m})}e^{-\frac{1}{2}\gamma|\bm{r}-\bm{r}_{\ell}|}{\,{\rm d}}\bm{r}\leq\frac{C_{d}}{\mathfrak{m}^{d}}\int_{\mathbb{R}^{d}}e^{-\frac{1}{2}\gamma|\bm{r}-\bm{r}_{\ell}|}{\,{\rm d}}\bm{r}$